Fluorines atomic number is 9 and its atomic mass is 19. How many neutrons does fluoride have?
Answer #1

Atomic number denotes the number of protons in an atom which is usually denoted by Z.

Here, Z = 9

Atomic mass is the total sum of protons and neutrons in an atom and is usually denoted by A.

Here, A = 19

So number of neutrons = A - Z = 19 - 9 = 10

So number of neutrons in fluorine is 10 neutrons